### v2

@@ Abstract
Genomic studies, including CRISPR-based PerturbSeq analyses, face a vast hypothesis space, while gene perturbations remain costly and time-consuming. Gene expression models based on graph neural networks are trained to predict the outcomes of gene perturbations to facilitate such experiments. Active learning methods are often employed to train these models due to the cost of the genomic experiments required to build the training set. However, poor model initialization in active learning can result in suboptimal early selections, wasting time and valuable resources. While typical active learning mitigates this issue over many iterations, the limited number of experimental cycles in genomic studies exacerbates the risk. To this end, we propose graph-based one-shot data selection methods for training gene expression models. Unlike active learning, one-shot data selection predefines the gene perturbations before training, hence removing the initialization bias. The data selection is motivated by theoretical studies of graph neural network generalization. The criteria are defined over the input graph and are optimized with submodular maximization. We compare them empirically to baselines and active learning methods that are state-of-the-art on this problem. The results demonstrate that graph-based one-shot data selection achieves comparable accuracy while alleviating the aforementioned risks.
@@
Active Learning Graph Neural Networks Gene Perturbation Prediction
@@ 1 Introduction
Genomic research enables the study of genetic factors underlying various health conditions, opening new avenues for therapeutic development. Techniques such as CRISPR interference (CRISPRi) and PerturbSeq [5] have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings. However, analyzing combinatorial gene perturbations remains challenging [38, 28], leading most studies to focus on single-gene perturbations. Moreover, most CRISPR-based PerturbSeq experiments are restricted to the order of a few hundreds of gene perturbations [10, 1, 31] due to budget and time constraints, despite the availability of over 20,000 potential gene targets. To complement experimental observations, several machine learning models have been developed to predict the outcome of gene perturbations on a given cell  [34, 7, 8, 13, 22, 23, 14].
@@
Figure 1: The difference between active learning and one-shot data-selection for genomic experiments with GEARS [34].
These models leverage diverse neural architectures, ranging from autoencoders [7] to graph neural networks [34], and they require a large number of training samples to generalize effectively. This reliance on supervision occasionally defeats their main purpose (i.e., minimizing the experimental cost), because the labels are themselves retrieved from genomic experiments. Additionally, disease-specific variations in gene expression make it difficult to train a universal model, necessitating the development of models tailored to specific experimental conditions.
Active learning [37] offers a promising solution by enabling iterative interaction between the wet lab and the machine learning models, to optimize data collection and hence training, with a limited number of cycles, as shown in Fig. 1(a). In the context of CRISPRi-based PerturbSeq experiments, these cycles tend to be restricted (e.g. 5 cycles), as each experiment in the lab may take up to a month [17]. Given the vast number of possible genes to explore, active learning strategies are well-suited to efficiently train predictive models while minimizing the experimental costs.
@@
However, active learning depends on the model’s initialization, which poses a non-negligible risk in this context. Fig. 2 illustrates the amount of variation induced in the genomic perturbations selected to train the GEARS model [34], based on 5 different random initializations. In particular, it illustrates the number of genes that are selected in any one given cycle as black bars and the total number of unique genes that are selected for the 5 differently initialized models as the orange bars. Since the orange bars are often more than twice as large than the black ones, we observe that the perturbed genes differ significantly for each initialization, i.e., an extreme sensitivity of the selected genes to the model’s initialization. The effect is more pronounced in the initial cycles, as the models become more stable in subsequent cycles and converge to similar choices. However, given the aforementioned limitation in the cycle horizon, the train set may mistakenly include gene perturbations that are difficult to correct afterwards. In typical active learning scenarios, a longer horizon leads to an overall stable selection, mitigating early-stage errors over time. Nevertheless, in this case, the wet lab may risk losing months of work and resources for a model that is not properly initialized and may not converge in time.
@@
Figure 2: The total number of unique genes perturbed in active learning experiments in two datasets with 5 different random seed initializations of GEARS [34], on the same train-test split. Compared to the budget for gene perturbations, i.e., the number of genes to be perturbed in each cycle, this indicates the variation associated with building the dataset based on the model in the initial cycles.
@@
To mitigate this risk, we explore the use of one-shot data selection methods as an alternative to active learning for the problem of training gene expression predictors. By selecting the samples upfront based on the input of the model, as seen in Fig. 1(b), we can provide competitive results while exhibiting:
@@
  • Independence of model initialization. A data-selection method runs before the model’s training and hence it does not depend on the model’s predictions. This reduces the risk of wasting experimental resources on erroneous gene perturbations.
@@
  • Reduced operational delay. Each iteration in active learning introduces extra computational overhead: model retraining, recalculating uncertainty scores, etc. Moreover iterative experiments exhibit delays due to logistical complexities and communication, compared to one large experiment. These problems can add a substantial delay, which is not present in one uniform experiment.
@@
We propose two one-shot data selection methods for training the state-of-the-art GEARS model [34], although our methods can work with any graph-based predictor. We rely on theorems for GNN generalization to define proxy criteria to guide the sample selection based on GEARS’ input knowledge graph. The criteria are optimized greedily to build the train set. Our contributions are as follows:
@@
  • Two novel data-selection criteria based on generalization theorems for GNNs that utilize the input knowledge graph and are model-free.
@@
  • A cost effective lazy forward submodular maximization algorithm to optimize one of the criteria, and a respective greedy heuristic for the other.
@@
  • An empirical comparison including two active learning methods (one of which is SOTA in our problem) and two baselines, indicating the effectiveness of one-shot data selection.
@@ 2 Related Work
Our one-shot data-selection methodology addresses this gap along with the fundamental problems of the active learning analyzed in sec. 1. We focus on training GEARS effectively while lacking prior information and the ability to take feedback from the model. The former is a plausible assumption since data for specific hypothesis are not that commonly available, and the latter is motivated by the variation inherent in the initial cycles of an active learning process and the cost to acquire samples in our context.
@@ 3 Methodology
In this section we first formulate the problem and clarify the motivation for pure graph-based methods for one-shot data selection. Subsequently we introduce the two methods for one-shot data selection.
@@ 3.1 Formulation
We are given a gene graph $G=(V,E),$ with $|V|=N$ and $|E|=M$, where some of the nodes are candidate perturbations. We assume that the dataset is partially labeled, meaning there is a subset of nodes $\mathcal{L}\subset V$ labeled. Each label is a matrix $Y_{u}\in\mathcal{R}^{c\times|g|}$ which represents the gene expression of a set of genes $g$ (which can belong to $V$ or not) of $c$ cells that underwent perturbation $u\in V$, along with the respective pre-perturbation gene expressions $X_{u}\in\mathcal{R}^{c\times g}$. There is respectively a set of genes $\𝑈\𝑉U=V\backslash\mathcal{L}$ that are unlabeled. GEARS learns how to predict $Y_{u}$ based on a given perturbation $u$. The problem we focus on is:
    Iteratively build the train set $S$ with the $b$ best genes to perturb from set $U$ to maximize GEARS’ accuracy on an unseen set of gene perturbations.
We aim to build the train set $S$ in order to cover as much and as diverse information as possible to train the model. Similar to the core-sets [36], we define a covering function that maximizes the model’s reach, but instead of defining it in the embedding space, we define it in the knowledge graph. The main reason behind this is that our graphs are not accompanied by node features. The node embeddings are randomly initialized and learned through training. This means any density-based method would be conditioned on the model initialization problem we explained in sec. 1.
To be specific, in GEARS and most gene expression models, the pre-perturbation gene expressions $X_{u}$ are not utilized during the prediction process as node features, but rather they are added to the predicted perturbation to arrive at $Y_{u}$. This happens because the prediction of the perturbation should be independent to the state of the cell $X_{u}$, as in real experiments, the cells will not exist until we have already decided which perturbation to run. This is why gene expressions models predict directly the gene expression displacement given the name of the perturbation, without using the cell $X_{u}$ [34, 7, 13]. Hence, in the absence of input node features and with the intention to remove the dependence of the train set selection on the model (due to the aforementioned variability in the resulting election), we are left with the knowledge graph as our sole remaining source of information to select the train set. The GEARS architecture is described in detail in the appendix for reference and further clarification.
@@ 3.2 GraphReach
To understand how to use the graph to guide our selection, we start with the graph convolution layer of GEARS. The Simplified Graph Convolution (SGC) layer [43] is defined as
$\tilde{\mathbf{A}}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ (1)
$\mathbf{H}_{t+1}=\tilde{\mathbf{A}}^{K}\mathbf{H}_{t}\mathbf{W}$ (2)
where $\mathbf{H}_{t}$ are the node representations at layer $t$, $\mathbf{W}$ represents the learnable parameters of the model, $\mathbf{A}$ is the graph’s adjacency matrix, $K$ is the power of the adjacency predefined as 1 in GEARS, and $\mathbf{D}$ is the diagonal degree matrix. It should be noted that SGC is known to exhibit better generalization gaps than most of the other layers [18, 41], rendering it a well-suited module for our task.
Intuitively, our main goal is to label nodes such that we cover as much information as possible from the input, akin to core-sets. As mentioned above, the SGC aggregates information from the whole network; yet only the representations of nodes that are within the reach of labeled training nodes will get updated through the message passing layers. For example, if there are $L$ layers, message passing reaches up to the $L$-hop neighborhood of labeled nodes in set $S$. These are the only nodes relevant to training, i.e., their embeddings are updated based on supervision signal, as the rest of the graph does not effectively contribute to the training process. It is important to highlight this semi-supervised setting allows for the update of test nodes’ representations if they are sufficiently close to the training nodes. This clarifies how covering bigger parts of the graph can potentially improve the prediction.
Thus, to maximize the reach of the supervision signal and cover as much of the input information as possible, we should aim to maximize the number of nodes reached by the training set $S$. Beyond our intuition, this is supported by theoretical insights into the generalization capability of GNNs. The number of nodes that we have to predict and the number of samples that we can add in each cycle is fixed, so we can not change the VC dimension [35]. However, theoretical studies suggest that the generalization of a GNN is improved based on the distance between the training and the test nodes in the graph [25, 46], supporting our intuition that covering a larger part of the graph with training nodes-and thereby reducing the distance to the test set- is beneficial to the model. Note that these works define the distance as the structural and distributional difference; nevertheless, since we lack node features, the structural distance is the sole factor in our case.
Based on the above, we can define our criterion function, similar to an acquisition function in active learning, such that we maximize the number of nodes reached by the set $S$. Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
@@
$R_{L}(S,A)=\{v\in V\mid(\mathbf{A}^{\ell}_{uv}>0\text{ for some }u\in S,\text{and }1\leq\ell\leq L\}$ (4)
The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments in GEARS. The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration. This is fundamentally suboptimal [6, 20] but many active learning methods [24, 30, 42, 40] prove the base function is monotonic and submodular to achieve a guarantee that the result will be at least $1-1/e$ close to the optimal [29]. We follow suit and prove the same for reachability.
@@
The greedy approach requires testing every available perturbation for every addition during a cycle, i.e., $\mathcal{O}(|U|T)$ where $T$ is the total training budget in samples. We employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy [21]. The final data selection algorithm, called GraphReach from graph reachability, is shown in Alg. 1.
@@ 4.1 Design
It should be noted that GEARS utilizes the whole graph to learn representations and uses the gene perturbation datasets for supervision only. In other words, the model uses the perturbations corresponding to the train set to compute the loss, but the SGC layer runs over the whole graph, including the test perturbations and unrelated genes, in a semi-supervised learning fashion [19]. The data selection methods follow suit and utilize the whole graph. For example the reachability of a gene perturbation in Replogle is not defined only on the 1081 perturbations, but on the 9,800 nodes.
@@ 4.2 Results
As can be seen in Tab. 1, the models developed from the data selection methods perform better than SOTA in one dataset and close to it in another. To be precise, while MaxSpec and TypiClust are the top methods in Reglogle and Adamson respectively, GraphReach has overall the most balanced performance being first or second in all datasets and metrics with minimal differences (considering the standard deviations) from the top.
Fig. 3(a) and 4(a) show that GraphReach outperforms marginally TypiClust in 3 out of 5 cycles in the Adamson dataset. Both achieve a better accuracy by the third cycle than the other models do in 5 cycles, with the exception of ACS-FW at the 5th cycle. The large standard deviations in Adamson stem from the reduced number of perturbations, which causes different test splits to exhibit significant variation in difficulty level. These variations are not directly associated with the variations from model initialization shown in Fig. 2, as they stem from the difference in data splits.
@@
Data Selection MaxSpec $\mathbf{150\pm 4}$ $160\pm 49$ $920\pm 9$ $958\pm 11$
@@
For the Replogle dataset, the active learning methods are outperformed clearly by the graph-based data selection methods and even by random selection in the initial cycles. ACS-FW in particular has considerably worse performance than random, which is why it is not included in Fig. 3(b) and Fig. 4(b), while TypiClust consistently falls behind the remaining methods. In contrast, MaxSpec is the best method in Replogle but argueably the worst in Adamson. We hypothesize that this performance difference is related the drastic difference in size of these two datasets with Replogle containing several hundreds of nodes, whereas Adamson only considers graphs composed of tens of nodes. GraphReach on the other hand, is the only consistent method in both datasets from all benchmarks. This happens because optimizing for reachability has the same effect independently of the graph’s scale: maximizing the coverage of the train set based on the input information.
@@
Regarding the course of the models’ accuracy throughout cycles, it is visible that adding wrong perturbations actually deteriorates the accuracy of the model instead of keeping it stable e.g. cycle 3 of TypiClust in Replogle or cycle 4 of ACS-FW in Adamson. This indicates that the base model can overfit significantly in case of wrong additions. In a realistic setting with limited gene perturbations for training and a much larger test set (e.g. all available genes), we expect this effect to be greater. This further supports the need for the data selection methods, which have no cycle of deteriorating accuracy.

### v3

@@ Abstract
Genomic studies, including CRISPR-based PerturbSeq analyses, face a vast hypothesis space, while gene perturbations remain costly and time-consuming. Gene expression models based on graph neural networks are trained to predict the outcomes of gene perturbations to facilitate such experiments. Active learning methods are often employed to train these models due to the cost of the genomic experiments required to build the training set. However, poor model initialization in active learning can result in suboptimal early selections, wasting time and valuable resources. While typical active learning mitigates this issue over many iterations, the limited number of experimental cycles in genomic studies exacerbates the risk. To this end, we propose graph-based data filtering as an alternative. Unlike active learning, data filtering selects the gene perturbations before training, meaning it is free of bias due to random initialization and initial random selection. Moreover, reducing the iterations between the wet lab and the model provides several operational advantages resulting in significant acceleration. The proposed methods are motivated by theoretical studies of graph neural network generalization. The criteria are defined over the input graph and are optimized with submodular maximization. We compare them empirically to baselines and active learning methods that are state-of-the-art. The results demonstrate that graph-based data filtering achieves comparable accuracy while alleviating the aforementioned risks.
@@
Gene Perturbation Prediction Active Learning Optimal Experimental Design
@@ 1 Introduction
Genomic research enables the study of genetic factors underlying various health conditions, opening new avenues for therapeutic development. Techniques such as CRISPR interference (CRISPRi) and PerturbSeq [5] have revolutionized the landscape of genomic experimentation by enabling high-throughput screenings. However, analyzing combinatorial gene perturbations remains challenging [39], leading most studies to focus on single-gene perturbations. Moreover, most CRISPR-based PerturbSeq experiments are restricted to the order of a few hundreds of gene perturbations [10, 1, 33] due to budget and time constraints, despite having over 20,000 potential gene targets. To assist exploring this vast space, machine learning models have been developed to predict the outcome of gene perturbations on a given cell  [36, 7, 8, 14, 25, 26, 15].
These models leverage diverse neural architectures, ranging from autoencoders [7] to graph neural networks [36], which require a substantial number of training samples to achieve generalization. These labeled samples stem from genomic experiments, which, as mentioned above, are rather costly. This, coupled with the fact that the models are disease specific, meaning they are trained on single cell data from the same hypotheses, render the model training an optimization problem in and of itself. Typically this is addressed with active learning [38] which defines an iterative interaction between the wet lab and the model in order to retrieve the genes that will optimize the training, as shown in Fig. 1(a). Active learning strategies can be based on i) uncertainty, which query instances where the model is least confident, ii) information theory, which seek instances that are expected to provide the most information about the model parameters and iii) density, which ensures that selected samples are representative of the input distribution. The first two strategies use the model trained on the available data to guide the selection at each iteration. On the contrary, density-based methods typically rely on the dataset distribution without input from the model, which, however, is not possible in the current setting. Specifically, in SOTA models such as GEARS [36], the only input available before executing the genomic experiment is the candidate gene perturbation name and a knowledge graph defined a priori. Hence, density-based methods rely on the model to map the genes to a learnable embedding space and then act upon it [18]. These renders all our available strategies model-based, which is highly problematic.
@@
Figure 1: The difference between active learning and data filtering for genomic experiments with GEARS [36].
[…]
@@
To quantify the sensitivity of the framework to initialization bias, Fig. 1 illustrates the amount of variation induced in the gene perturbations selected to train GEARS, based on 5 different random initializations of the model using the SOTA algorithm for the problem [18]. In particular, it illustrates the number of genes that are selected in any one given cycle as black bars and the total number of unique genes that are selected for the 5 differently initialized models as the orange bars. Since the orange bars are often more than twice as large as the black ones, we observe that the perturbed genes differ significantly for each initialization, i.e., an extreme sensitivity of the selected genes to the model’s initialization. The effect is more pronounced in the initial cycles because the models become more stable in subsequent cycles and converge to similar choices, as expected by an AL procedure. However, given the aforementioned limitation in the cycle horizon, the train set does not recover.
@@
Figure 2: The total number of unique genes perturbed in active learning experiments in two datasets with 5 different random seed initializations of GEARS [36], on the same train-test split. Compared to the budget for gene perturbations, i.e., the number of genes to be perturbed in each cycle, this indicates the sensitivity of the gene selection process to the model initialization.
@@
To mitigate this risk, we explore methodologies that disentangle the model with the selection strategy. Specifically, we focus on data filtering methods as an alternative to active learning [22]. Data filtering selects samples before training the model, based on the information contained a priori. By selecting the samples upfront, as seen in Fig. 1(b), we can provide competitive results while exhibiting several key advantages:
@@
  • Independence of model initialization. Data-filtering does not depend on the model’s predictions. This reduces the risk of wasting experimental resources on erroneous gene perturbations due to lack of training data, random model initialization and reduced iteration number.
  • Decrese experimental time. In active learning, each iteration requires the wet lab to run experiments, which may last weeks each. In contrast, with data-filtering, we ask for the same number of experiments but only once. Since the cell assays can run in parallel, this means data filtering will be solved roughly in the time required for one iteration of active learning. This translates to 3-4 months of acceleration in real life.
@@
  • Reduced operational delay. Each iteration requires a computational overhead for model retraining, computing the acquisition step, etc. Moreover, iterative experiments exhibit non-negligible delays due to logistical complexities and communication between ML team and the wetlab. These problems can add a substantial delay, which is not present when we only run one experiment.
@@
We propose two data filtering methods for training the GEARS model, although our methods can work with any graph-based predictor. We rely on theorems for GNN generalization to define proxy criteria to guide the sample selection based on GEARS’ input knowledge graph. The criteria are optimized greedily to build the training set. Note that, we focus solely on the knowledge graph because the coexpression network is derived from the cell assays, which by default are not available before the experiment takes place.
Our contributions are as follows:
@@
  • Two novel data-filtering criteria based on generalization theorems for GNNs that utilize the input knowledge graph and are model-free.
@@
  • A submodular maximization algorithm to optimize one of the criteria, and a respective greedy heuristic for the other.
@@
  • An empirical comparison including two active learning methods (one of which is SOTA in our problem) and two baselines, indicating the effectiveness of data filtering.
@@ 2 Related Work
Our data filtering methodology addresses this gap along with the fundamental problems of the active learning analyzed in sec. 1. We focus on training GEARS effectively while lacking prior information and the ability to take feedback from the model. The former is a plausible assumption since data for specific hypotheses are not that commonly available, and the latter is motivated by the variation inherent in the initial cycles of an active learning process and the cost to acquire samples in our context.
@@ 3 Methodology
In this section, we first formulate the problem and clarify the motivation for graph-based approaches, and subsequently we introduce the two methods for data filtering.
@@ 3.1 Formulation
We are given a gene knowledge graph $G=(V,E)$, where $|V|=N$ and $|E|=M$, and each node represents a gene, some of which correspond to candidate gene perturbations. Our goal is to select a subset $S\subset V$, under a budget constraint $|S|=b$ with $b\ll N$, to experimentally perturb in a wet lab setting.
For each selected gene $u\in S$, the perturbation experiment yields paired measurements of pre- and post-perturbation gene expression: $X\in\mathbb{R}^{c\times g}$ for $c$ single-cell profiles and $g$ expressed genes (not to be confused with the perturbed genes), and corresponding post-perturbation outputs $Y\in\mathbb{R}^{c\times g}$. Thus, the perturbation effect is modeled as an additive change $\hat{P}_{u}\in\mathbb{R}^{g}$ such that $\hat{Y}_{u}=X+\hat{P}_{u}$, and the model learns a mapping $M:V\rightarrow\mathbb{R}^{g}$. Note that the perturbation is run on multiple cells with varying gene expressions.
The model $\theta^{S}$, trained only on the selected set $S$, is a graph neural network (GNN) such as GEARS. The loss function is the mean squared error between predicted and observed expression:
$\ell(X,Y,u;\theta^{S})=\|Y-(X+\hat{P}_{u})\|^{2},$ (1)
where $\hat{P}_{u}=M(u;\theta^{S})$. The problem we address is to select the most informative subset of genes to perturb, such that the trained model generalizes well to unseen perturbations $U\subseteq V\setminus S$. Formally, we aim to:
$\underset{S\subset V,\ |S|=b}{\arg\min}\ \sum_{u\in U}\ell(X_{u},Y_{u},u;\theta^{S})\leavevmode\nobreak\ ,\leavevmode\nobreak\ s.t.|S|=b,$ (2)
where $X_{u},Y_{u}$ denote the inputs and targets for perturbation $u$, and $\theta^{S}$ is trained only on data from $S$.
3.2 Motivation
As mentioned above, in normal density-based active learning we try to cover as much of the input space as possible. For example, in core-sets[37], it is proven for i.i.d. data that if the training samples are selected such that each test sample is less then $\delta$ away from one of them, we are guaranteed to reduce generalization error with a factor analogous to $\delta$ and the total number of samples. This can be seen as ’covering’ the covariate space with the training samples that have a $\delta$-radius. Their main assumption is that samples that are close in space exhibit similar outcomes. In our case this assumption translates to homophily ,i.e. , the effects of genes that are close in the KG are similar, and the $\delta$ radius can be defined based on hop distance. However, since GNNs utilize message passing, they break the i.i.d. assumptions which makes this result unusable.
In terms of active learning for GNNs, labeling a node incurs an information gain to every other node in their receptive field [43, 46, 45] (,i.e, $k$-hop neighborhood), meaning that the supervision signal propagates due to message passing. In the same vain, [19] argue that training is improved by maximizing the diversity of aggregated features, which translates to maximizing the number of unique nodes in the training. The problem with these approaches is that the metrics depend on the node distances, the features of unlabeled nodes, and the representations of the model. In our setting, $X$ and $Y$ are available only for labeled nodes, and we have no access to the model for the reasons detailed in section 1, hence we can not directly utilize this framework.
These approaches agree with the basic intuition that although a GNN aggregates information from the whole network, the only representations that will get updated correspond to the nodes that are within the reach of the training nodes. For example, if there are $L$ layers, message passing reaches up to the $L$-hop neighborhood of labeled nodes in set $S$. These are the only nodes relevant to training, i.e., their embeddings are updated based on supervision signal, as the rest of the graph does not effectively contribute to the training process. It is important to highlight this semi-supervised setting allows for the update of test nodes’ representations if they are sufficiently close to the training nodes. This clarifies how covering bigger parts of the graph can potentially improve the prediction.
Thus, to maximize the reach of the supervision signal and cover as much of the input information as possible, we should aim to maximize the number of nodes reached by the training set $S$. Beyond the aforementioned literature, this is supported by theoretical insights into the generalization capability of GNNs. Theoretical studies suggest that the generalization is improved as we reduce the distance between the training and the test nodes in the graph [28, 48]. This supports our intuition that covering a larger part of the graph with training nodes-and thereby reducing the distance to the test set- is beneficial to the model. Note that these works define the distance as the structural and distributional difference; nevertheless, since we do not have access to the features or the model, we will define this covering functions on the only input space we have, the knowledge graph.
Figure 3: An overview of the data filtering methodology. Our sole input is the knowledge graph from GO[36] and a list of candidate perturbations. The data filtering algorithm ,e.g. , GraphReach selects the set of gene perturbations and they are given to the wet lab for the experimental part. Finally, the single cell gene expression data is given to GEARS for training and validation.
@@ 3.3 GraphReach
We define our criterion function, similar to an acquisition function in active learning, such that we maximize the number of nodes reached by the set $S$. Specifically for a new node $v$, the criterion is defined as the additional reachability that $v$ adds in the set of train perturbations $S$
@@
$R_{L}(S,A)=\{v\in V\mid\mathbf{A}^{\ell}_{uv}>0\text{ for some }u\in S,\text{ and }1\leq\ell\leq L\}$ (4)
The subscript $L$ is omitted in subsequent discussion as it remains stable throughout all experiments. The criterion function selects greedily the node $v$ maximizing $\alpha(v,S)$ in every iteration. This is fundamentally suboptimal [6, 23] but many active learning methods [27, 32, 42, 41] prove the base function is monotonic and submodular to achieve a guarantee that the result will be at least $1-1/e$ close to the optimal [31]. We follow suit and prove the same for reachability.
@@
The greedy approach requires testing every available perturbation for every addition during a cycle, i.e., $\mathcal{O}(|U|T)$ where $T$ is the total training budget in samples. We employ a cost effective lazy forward approach to accelerate the acquisition step without loss of accuracy [24]. The final data filtering algorithm, called GraphReach from graph reachability, is shown in Alg. 1, and an overview of the overall step-by-step data filtering approach can be seen in Fig. 3.
@@ 4.1 Design
It should be noted that GEARS utilizes the whole knowledge graph to learn representations and uses the gene perturbation datasets for supervision only. In other words, the model uses the perturbations corresponding to the train set to compute the loss, but the SGC layer runs over the whole graph, including the test perturbations and unrelated genes, in a semi-supervised learning fashion [21]. The data filtering methods follow suit and utilize the whole graph. For example, the reachability of a gene perturbation in Replogle is not defined only on the 1088 candidate perturbations, but on the whole 9,800 nodes.
@@ 4.2 Results
As can be seen in Tab. 1, the models developed from the data filtering methods perform better than SOTA in one dataset and close to it in another. To be precise, while MaxSpec and TypiClust are the top methods in Reglogle and Adamson respectively, GraphReach has overall the most balanced performance being first or second in all datasets and metrics with minimal differences (considering the standard deviations) from the top.
Fig. 4(a) and 5(a) show that GraphReach outperforms marginally TypiClust in 3 out of 5 cycles in the Adamson dataset. Both achieve a better accuracy by the third cycle than the other models do in 5 cycles, with the exception of ACS-FW at the 5th cycle. The large standard deviations in Adamson stem from the reduced number of perturbations, which causes different test splits to exhibit significant variation in difficulty level. These variations are not directly associated with the variations from model initialization shown in Fig. 2, as they stem from the difference in data splits.
@@
Data Filtering MaxSpec $\mathbf{150\pm 4}$ $160\pm 49$ $920\pm 9$ $958\pm 11$
@@
For the Replogle dataset, the active learning methods are outperformed clearly by the graph-based data filtering methods and even by random selection in the initial cycles. ACS-FW in particular has considerably worse performance than random, which is why it is not included in Fig. 4(b) and Fig. 5(b), while TypiClust consistently falls behind the remaining methods. In contrast, MaxSpec is the best method in Replogle but argueably the worst in Adamson. We hypothesize that this performance difference is related the drastic difference in size of these two datasets with Replogle containing several hundreds of nodes, whereas Adamson only considers graphs composed of tens of nodes. GraphReach on the other hand, is the only consistent method in both datasets from all benchmarks. This happens because optimizing for reachability has the same effect independently of the graph’s scale: maximizing the coverage of the train set based on the input information.
@@
Regarding the course of the models’ accuracy throughout cycles, it is visible that adding wrong perturbations actually deteriorates the accuracy of the model instead of keeping it stable e.g. cycle 3 of TypiClust in Replogle or cycle 4 of ACS-FW in Adamson. This indicates that the base model can overfit significantly in case of wrong additions. In a realistic setting with limited gene perturbations for training and a much larger test set (e.g. all available genes), we expect this effect to be greater. This further supports the need for the data filtering methods, which have no cycle of deteriorating accuracy.